Jessica has half of her investments in stock paying a
7%
dividend and the other half in a stock paying
13%
interest. If her total annual interest is
$780
, how much does she have invested?

Found 2 solutions by richwmiller, lwsshak3:
Answer by richwmiller(17219)

Stop spreading the problem out to make it look bigger.
We know that the interest for the two accounts is $780
0.07*x+0.13*y=780
We know that the account at 7% has 1 times the amount at 13%
We substitute for x
0.07*(1*y)+0.13*y=780
We multiply out
(0.07+0.13)*y=780
We combine like terms.
0.2*y=780
Isolate y
y= 3900.00
y=$ 3900.00 at 13%
Calculate x
x=1*y
x=$ 3900.00 at 7%
Your total for x is $ 3900.00
Your total for y is $ 3900.00
Your total for x and y is $ 7800.00
Total invested $ 3900.00 + $ 3900.00=$ 7800.00
We check
0.07* 3900.00+0.13* 3900.00=780
273.00+ 507.00=780
780=780
Since this statement is TRUE and neither x nor y is negative then all is well.
codeintry

Answer by lwsshak3(11628)

Jessica has half of her investments in stock paying a
7% dividend and the other half in a stock paying13%
interest. If her total annual interest is $780,how much does she have invested?
***
let x=amt Jessica invested
7%*x/2+13%*x/2=780
.07x+.13x=1560
.20x=1560
x=7800
how much does she have invested? $7800